# All-atom Diffusion Transformers: Unified generative modelling of molecules and materials

### v1

@@ Abstract
Diffusion models are the standard toolkit for generative modelling of 3D atomic systems. However, for different types of atomic systems – such as molecules and materials – the generative processes are usually highly specific to the target system despite the underlying physics being the same. We introduce the All-atom Diffusion Transformer (ADiT), a unified latent diffusion framework for jointly generating both periodic materials and non-periodic molecular systems using the same model: (1) An autoencoder maps a unified, all-atom representations of molecules and materials to a shared latent embedding space; and (2) A diffusion model is trained to generate new latent embeddings that the autoencoder can decode to sample new molecules or materials. Experiments on QM9 and MP20 datasets demonstrate that jointly trained ADiT generates realistic and valid molecules as well as materials, exceeding state-of-the-art results from molecule and crystal-specific models. ADiT uses standard Transformers for both the autoencoder and diffusion model, resulting in significant speedups during training and inference compared to equivariant diffusion models. Scaling ADiT up to half a billion parameters predictably improves performance, representing a step towards broadly generalizable foundation models for generative chemistry.
@@ 1 Introduction
All atomic systems share the same underlying physical principles that determine their 3D structure and interactions. However, we currently do not have a unified formulation of diffusion models across different types of atomic systems such as small molecules, biomolecules, crystals, and their combinations. Most diffusion models are highly specific to each type of system, and involve multi-modal generative processes on complex product manifolds of categorical and continuous data types. For example, de novo generation of small molecules is modelled as two independent diffusion processes for the atom types (categorical) and 3D coordinates (continuous) of a set of atoms (Hoogeboom et al., 2022). The denoiser model learns how atom types and 3D coordinates jointly evolve in order to sample new molecules but passes through unrealistic intermediate states during the denoising trajectory. Diffusion models for biomolecules treat groups of atoms as rigid bodies and add a third manifold (rotations) into the joint diffusion process (Yim et al., 2023b; Campbell et al., 2024). For crystals/materials, the diffusion process needs to additionally handle periodicity and operates on a joint manifold of atom types, fractional coordinates, lattice lengths, and lattice angles that together define the repeating unit cell (Xie et al., 2022; Miller et al., 2024).
@@
ADiTs can be trained jointly on both periodic and non-periodic 3D atomic systems, demonstrating broad generalizability. Training a single unified model on the QM9 molecular and MP20 materials datasets leads to state-of-the-art performance in both domains, exceeding specialized equivariant diffusion models on physics-based validations. DFT calculations reveal that ADiTs generate stable, unique, and novel crystals at a 5-6% S.U.N. rate, a 25% improvement upon the 4-5% rates of previous methods. Joint training yields higher validity rates than QM9-only or MP20-only ADiT variants, demonstrating successful transfer learning between periodic and non-periodic atomic systems.
@@ 3 Experimental Setup
Datasets. We train models on periodic crystals from MP20 and non-periodic molecules from QM9, representing two distinct domains of atomic systems. MP20 (Xie et al., 2022) contains 45,231 metastable crystal structures from the Materials Project (Jain et al., 2013), each with up to 20 atoms and spanning 89 different element types. QM9 (Wu et al., 2018) consists of 130,000 stable small organic molecules containing up to nine heavy atoms (C, N, O, F) along with hydrogens. We split the data following prior work (Xie et al., 2022; Hoogeboom et al., 2022) to ensure fair comparisons.
@@ 4 Results
Table 2: Molecule generation results. We report (a) validity and uniqueness rates, as well as (b) % pass rates on 7 sanity checks from Posebusters for 10,000 sampled molecules. ADiTs match or improve performance w.r.t. baselines, and sample physically realistic structures. Joint training with both molecular and crystal data improves molecular generation performance over QM9-only models. ( denotes models which explicitly generate hydrogen atoms.)
@@
Figure 2: ADiTs are significantly faster than equivariant diffusion models. We plot the validity rate vs. number of integration steps for ADiTs and equivariant diffusion models for generating 10,000 samples on a single V100 GPU. ADiTs scale significantly better with the number of integration steps compared to equivariant diffusion.
@@
Extension to MOF generation. Having established that ADiTs benefit from transfer learning to generate high-quality crystals and molecules, we next challenged our models to generate metal-organic frameworks (MOFs), which represent a more complex class of hybrid materials with metal nodes connected by organic small molecule linkers. We trained ADiTs on an additional 14,000 MOFs of up to 150 atoms from the QMOF database (Rosen et al., 2021) alongside QM9 and MP20, using the same experimental settings and training for 10,000 epochs. We sampled 1,000 MOFs and evaluated their validity using 15 sanity checks from MOFChecker (Jablonka, 2023), including tests for: presence of metal/carbon/hydrogen atoms, atomic overlaps, over/undervalent carbons and nitrogens, missing hydrogens, and excessive partial charges.
Table 3 shows that QMOF-only trained ADiT achieves a 15% overall validity rate for MOF generation, which decreases to 10% with joint training on molecules, crystals, and MOFs simultaneously. However, the jointly trained model takes significant time to train and did not fully converge, suggesting that further improvements in MOF generation may be possible with larger models trained for longer. Comparing validity rates for joint vs. dataset-specific ADiTs shows that the joint model benefits from transfer learning across molecules, crystals, and MOFs, achieving high validity rates earlier in training. Notably, the joint model achieves high validity rates of 91% for crystals and 95% for molecules, matching our best dataset-specific models and being able to additionally generate MOFs. While not directly comparable, models specialized for MOF such as MOFDiff (Fu et al., 2024) are trained on large synthetic dataset of 300,000 MOFs and achieve 30% validity rates based on MOFChecker after DFT relaxation of generated MOFs (our results are reported without DFT relaxation).
@@
Additional results and visualizations. For additional results and ablation studies on ADiT, see Appendix C and Appendix D, respectively. In Appendix E, we further analyze ADiT’s joint latent space via PCA and visualize some sampled crystals, molecules, and MOFs.
@@
5 Discussions
Our work represents a significant step towards a broadly applicable foundation models for generative chemistry. We have introduced a unified latent diffusion framework for generating molecules and materials using a single model, and demonstrated the benefits of transfer learning from diverse atomic systems. The All-atom Diffusion Transformer (ADiT) matches or exceeds state-of-the-art performance while being conceptually simpler and computationally more efficient than previous domain-specific approaches.
However, several limitations point to promising future directions. First, we currently use relatively small datasets for training, which may limit model generalization. Scaling to larger and more diverse datasets such as Alexandria and the Cambridge Structural Database for crystals, ZINC for small molecules, and the Protein Data Bank for biomolecular complexes could significantly improve performance and enable learning of broadly applicable chemical principles. Second, while we demonstrate success on small molecules and crystals of up to tens of atoms, we have not yet fully validated our approach on larger systems such as metal-organic frameworks or biomolecules containing thousands of atoms, though initial results for MOF generation are promising. Recent work on biomolecular structure prediction with AlphaFold3 (Abramson et al., 2024) demonstrates that simple Gaussian diffusion models with standard Transformers can effectively handle systems with thousands of atoms. Adapting ADiT to larger scales, while maintaining its unified representation across periodic and non-periodic systems, could enable powerful transfer learning capabilities - especially valuable for low-data domains. Finally, our current models only perform unconditional generation – extending to conditional generation based on experimental properties, motif scaffolding, or molecular infilling would enable practical inverse design applications in drug discovery, materials science, and beyond.
Appendix A Related Work
Generative models for molecules and materials. Diffusion models have emerged as the state-of-the-art for generative modelling of atomic systems, with applications to molecules, crystals, and biomolecules. For molecule generation, Equivariant Diffusion (Hoogeboom et al., 2022) pioneered roto-translationally equivariant diffusion on the multi-modal product manifold of atom types and 3D positions, while GeoLDM (Xu et al., 2023) introduced latent diffusion in the space of an equivariant autoencoder. Schneuing et al. (2024) extended equivariant diffusion to generate molecules conditioned on binding protein partners for structure-based drug design, while Corso et al. (2023) explored similar architectures for protein-small molecule docking. For crystal generation, state-of-the-art approaches use equivariant diffusion on product manifolds of atom types, 3D/fractional coordinates, and lattice parameters. Notable examples include CDVAE (Xie et al., 2022), DiffCSP (Jiao et al., 2023), and FlowMM (Miller et al., 2024). MatterGen (Zeni et al., 2025) demonstrated conditional diffusion for inverse design based on target material properties and symmetry space groups.
Our work stands out as the first to develop unified generative models for both periodic and non-periodic atomic systems using latent diffusion. The closest work to ADiT in terms of diffusion formulation is AlphaFold3 (Abramson et al., 2024), which applies standard Transformers and Gaussian diffusion to generate all-atom biomolecular complex. However, their formulation is specific to structure prediction for biomolecules and only diffuses 3D atomic coordinates in Cartesian space. In contrast, our latent diffusion formulation is sufficiently general to work with both periodic and non-periodic systems, generating atom types, coordinates, as well as unit cell parameters unconditionally or with classifier-free guidance. Our emphasis on joint representations of molecules and crystals also aligns with recent work on general-purpose foundation models for molecular dynamics (Shoghi et al., 2024; Batatia et al., 2023). Similarly, our unified latent diffusion framework can potentially be scaled up with larger and more diverse chemical datasets towards foundation models for generative chemistry.
Latent diffusion models. Latent diffusion models (Vahdat et al., 2021; Rombach et al., 2022) propose to do diffusion in the latent space of an autoencoder instead of the raw input space of high-dimensional continuous signals such as pixels, and have been extremely successful for generating images, audio, and videos (Esser et al., 2024; Betker et al., 2023; Brooks et al., 2024). Latent diffusion is a more computationally efficient alternative to standard diffusion as the autoencoder’s latent space captures semantically meaningful features of the data, allowing for more efficient diffusion in a lower-dimensional space followed by reconstruction to the original data space. The original formulation was further improved by Diffusion Transformers (DiTs) (Peebles and Xie, 2023), which demonstrated that standard Transformers provide a highly scalable architecture for the denoiser network. Recent work by Ma et al. (2024) showed that DiT performance can be further enhanced by incorporating ideas from flow matching. Latent diffusion models can easily incorporate conditioning on additional information like class labels, text prompts, or infilling masks through classifier-based (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2022) as well as finetuning (Zhang et al., 2023; Dai et al., 2023).
However, equivariant networks are computationally expensive and harder to scale than standard Transformers in terms of data and model size. This is especially relevant for diffusion models, where denoisers typically process inputs as fully connected graphs to capture global structure (Joshi, 2020) and are iteratively run hundreds of times during inference. Recent work has challenges the necessity of equivariance for generative structure prediction tasks, showing that standard Transformers can achieve strong performance on biomolecular complexes (Abramson et al., 2024) and small molecule conformations (Wang et al., 2024). Non-equivariant models have also shown promising results for protein structure generation (Chu et al., 2024; Martinkus et al., 2024). In the same vein, our work leverages the simplicity and scalability of standard Transformers for generative modelling across both periodic and non-periodic 3D atomic systems, demonstrating that explicit equivariance is not a strict requirement for generating valid and realistic atomic structures at scale.
Appendix B Evaluation Metrics
  • Stability: % of crystals with DFT energy above hull $<$0.0 eV/atom and number of unique elements >=>=> = 2. (We also report metastability as DFT energy above hull $<$0.1 eV/atom and number of unique elements >=>=> = 2.)
  • Validity: % of molecules with canonical SMILES string found by RDKit.
  • Uniqueness: % of unique SMILES among valid ones.
Appendix C Additional Results

### v2

@@ Abstract
Diffusion models are the standard toolkit for generative modelling of 3D atomic systems. However, for different types of atomic systems – such as molecules and materials – the generative processes are usually highly specific to the target system despite the underlying physics being the same. We introduce the All-atom Diffusion Transformer (ADiT), a unified latent diffusion framework for jointly generating both periodic materials and non-periodic molecular systems using the same model: (1) An autoencoder maps a unified, all-atom representations of molecules and materials to a shared latent embedding space; and (2) A diffusion model is trained to generate new latent embeddings that the autoencoder can decode to sample new molecules or materials. Experiments on MP20, QM9 and GEOM-DRUGS datasets demonstrate that jointly trained ADiT generates realistic and valid molecules as well as materials, obtaining state-of-the-art results on par with molecule and crystal-specific models. ADiT uses standard Transformers with minimal inductive biases for both the autoencoder and diffusion model, resulting in significant speedups during training and inference compared to equivariant diffusion models. Scaling ADiT up to half a billion parameters predictably improves performance, representing a step towards broadly generalizable foundation models for generative chemistry.
@@ 1 Introduction
All atomic systems share the same underlying physical principles that determine their 3D structure and interactions. However, we currently do not have a unified formulation of diffusion models across different types of atomic systems such as small molecules, biomolecules, crystals, and their combinations. Most diffusion models are highly specific to each type of system, and involve multi-modal generative processes on complex product manifolds of categorical and continuous data types. For example, de novo generation of small molecules is modelled as two independent diffusion processes for the atom types (categorical) and 3D coordinates (continuous) of a set of atoms (Hoogeboom et al., 2022). The denoiser model learns how atom types and 3D coordinates jointly evolve in order to sample new molecules but passes through unrealistic intermediate states during the denoising trajectory. Diffusion models for biomolecules treat groups of atoms as rigid bodies and add a third manifold (rotations) into the joint diffusion process (Yim et al., 2023b; Campbell et al., 2024). For crystals and materials, the diffusion process needs to additionally handle periodicity and operates on a joint manifold of atom types, fractional coordinates, lattice lengths, and lattice angles that together define the repeating unit cell (Xie et al., 2022; Miller et al., 2024).
@@
ADiTs can be trained jointly on both periodic and non-periodic atomic systems, demonstrating broad generalizability. Training a single unified model on the QM9 molecular and MP20 materials datasets leads to state-of-the-art performance in both domains, exceeding specialized equivariant diffusion models on physics-based validations. DFT calculations reveal that ADiTs generate stable, unique, and novel crystals at a 5-6% S.U.N. rate, a 25% improvement upon the 4-5% rates of previous methods. Joint training yields higher validity rates than QM9-only or MP20-only ADiT variants, demonstrating successful transfer learning between periodic and non-periodic systems. ADiTs also match or exceed state-of-the-art equivariant models on the GEOM-DRUGS dataset of molecules with hundreds of atoms.
@@ 3 Experimental Setup
Datasets. For our main experiments, we train models on periodic crystals from MP20 and non-periodic molecules from QM9, representing two distinct domains of atomic systems. MP20 (Xie et al., 2022) contains 45,231 metastable crystal structures from the Materials Project (Jain et al., 2013), each with up to 20 atoms in its unit cell and spanning 89 different element types. QM9 (Wu et al., 2018) consists of 130,000 stable small organic molecules containing up to nine heavy atoms (C, N, O, F) along with hydrogens. We split the data following prior work (Xie et al., 2022; Hoogeboom et al., 2022) to ensure fair comparisons. We also include results on the GEOM-DRUGS dataset of 430,000 large organic molecules up to 180 atoms (Axelrod and Gomez-Bombarelli, 2022), as well as the QMOF dataset of 14,000 metal-organic framework structures (Rosen et al., 2021).
@@ 4 Results
Table 2: Molecule generation results on QM9. We report (a) validity and uniqueness rates, as well as (b) % pass rates on 7 sanity checks from Posebusters for 10,000 sampled molecules. ADiTs match or improve performance w.r.t. baselines, and sample physically realistic structures. Joint training with both molecular and crystal data improves molecular generation performance over QM9-only models. ( denotes models which explicitly generate hydrogen atoms.)
@@
Figure 2: ADiTs are significantly faster than equivariant diffusion models. We plot the number of integration steps for ADiTs and equivariant diffusion models vs. time to generate 10,000 samples on a single V100 GPU. ADiTs scale significantly better with the number of integration steps compared to equivariant diffusion.
@@
Extension to larger GEOM-DRUGS molecules. To demonstrate the scalability of the ADiT architecture to larger systems, we experiment with the GEOM-DRUGS dataset of 430,000 molecules of up to 180 atoms. Our setup follows Vignac et al. (2023) and we compare to state-of-the-art equivariant diffusion (Le et al., 2024) and flow matching (Irwin et al., 2025) baselines. In Table 3, we see that ADiT is on par or better than equivariant models across validity and PoseBusters metrics (Buttenschoen et al., 2025). This is a notable result because ADiT is based on the standard Transformer architecture with minimal molecular inductive biases and, unlike equivariant baselines, does not explicitly predict atomic bonds.
Table 3: Molecule generation results on GEOM-DRUGS. Left: Validity, uniqueness and % pass rates on Posebusters for 10,000 sampled molecules ( PoseBusters results taken from Buttenschoen et al. (2025)). ADiT with minimal molecular inductive biases matches or exceeds state-of-the-art equivariant diffusion baselines, which explicitly predict atomic bonds. Right: We plot the number of integration steps for ADiTs and SemlaFlow vs. time to generate 10,000 molecules on a single A100 GPU. ADiTs scale favorably with the number of integration steps compared to SemlaFlow, a highly optimized equivariant diffusion model.
Metric (pass) $\uparrow$ EQGAT-diff SemlaFlow ADiT
Validity 94.6 93.9 95.3
Uniqueness 100.0 100.0 100.0
Atoms connected 84.4 92.3 93.0
Bond angles 86.9 94.8 92.3
Bond lengths 87.0 94.6 92.5
Ring flat 87.0 94.9 95.4
Double bond flat 87.0 94.2 95.3
Internal energy 86.8 94.8 91.3
No steric clash 82.9 92.0 91.8
PoseBusters valid 59.7 87.5 85.3
@@
Additional results, ablations, and visualizations. For additional results and ablation studies, including extensions to metal-organic framework (MOF) generation, see Appendix C and Appendix D, respectively. In Appendix E, we further analyze transfer learning in ADiT’s joint latent space via PCA and visualize some sampled crystals, molecules, and MOFs.
5 Discussions
Our work represents a significant step towards a broadly applicable foundation model for generative chemistry. We have introduced a unified latent diffusion framework for generating molecules and materials using a single model, and demonstrated the benefits of transfer learning from diverse atomic systems. The All-atom Diffusion Transformer (ADiT) obtains state-of-the-art performance in molecular generation using an architecture based primarily on standard Transformers with minimal inductive biases. This makes ADiT conceptually simpler and computationally more efficient than previous domain-specific approaches based on equivariant diffusion.
However, several limitations point to promising future directions. First, we currently use relatively small datasets for training, which may limit model generalization. Scaling to larger and more diverse datasets such as Alexandria and the Cambridge Structural Database for crystals, ZINC for small molecules, and the Protein Data Bank for biomolecular complexes could significantly improve performance and enable learning of broadly applicable chemical principles. Second, while we demonstrate success on small molecules and crystals of up to hundreds of atoms, we have not yet fully validated our approach on larger systems such as metal-organic frameworks or biomolecules containing thousands of atoms, though initial results for MOF generation are promising. Recent work on biomolecular structure prediction with AlphaFold3 (Abramson et al., 2024) demonstrates that simple Gaussian diffusion models with standard Transformers can effectively handle systems with thousands of atoms. Adapting ADiT to larger scales, while maintaining its unified representation across periodic and non-periodic systems, could enable powerful transfer learning capabilities – especially valuable for low-data domains. Finally, our current models only perform unconditional generation – extending to conditional generation based on experimental properties, motif scaffolding, or molecular infilling would enable practical inverse design applications in drug discovery, materials science, and beyond.
Appendix A Related Work
Generative models for molecules and materials. Diffusion models have emerged as the state-of-the-art for generative modelling of atomic systems, with applications to molecules, crystals, and biomolecules. For molecule generation, Equivariant Diffusion (Hoogeboom et al., 2022) pioneered roto-translationally equivariant diffusion on the multi-modal product manifold of atom types and 3D positions, while GeoLDM (Xu et al., 2023) introduced latent diffusion in the space of an equivariant autoencoder. Schneuing et al. (2024) extended equivariant diffusion to generate molecules conditioned on binding protein partners for structure-based drug design, while Corso et al. (2023) explored similar architectures for protein-small molecule docking. For crystal generation, state-of-the-art approaches use equivariant diffusion on product manifolds of atom types, 3D/fractional coordinates, and lattice parameters. Notable examples include CDVAE (Xie et al., 2022), DiffCSP (Jiao et al., 2023), and FlowMM (Miller et al., 2024). MatterGen (Zeni et al., 2025) demonstrated conditional diffusion for inverse design based on target material properties and symmetry space groups. Language models have also been used for generating molecules and crystals as textual representations (Flam-Shepherd and Aspuru-Guzik, 2023; Gruver et al., 2024).
Our work stands out as the first to develop unified generative models capable of sampling both periodic crystals and non-periodic molecular systems jointly. The closest work to ADiT in terms of diffusion formulation is AlphaFold3 (Abramson et al., 2024), which applies standard Transformers and Gaussian diffusion to generate all-atom biomolecular complex. However, their formulation is specific to structure prediction for biomolecules and only diffuses 3D atomic coordinates in Cartesian space. In contrast, our latent diffusion formulation is sufficiently general to work with both periodic and non-periodic systems, generating atom types, coordinates, as well as unit cell parameters unconditionally or with classifier-free guidance. Our emphasis on joint representations of molecules and crystals also aligns with recent work on general-purpose foundation models for molecular dynamics (Shoghi et al., 2024; Batatia et al., 2023). Similarly, our unified latent diffusion framework can potentially be scaled up with larger and more diverse chemical datasets towards foundation models for generative chemistry.
Latent diffusion models. Latent diffusion models (Vahdat et al., 2021; Rombach et al., 2022) propose to do diffusion in the latent space of an autoencoder instead of the raw input space of high-dimensional continuous signals such as pixels, and have been extremely successful for generating images, audio, and videos (Esser et al., 2024; Betker et al., 2023; Brooks et al., 2024). Latent diffusion is a more computationally efficient alternative to standard diffusion as the autoencoder’s latent space captures semantically meaningful features of the data, allowing for more efficient diffusion in a lower-dimensional space followed by reconstruction to the original data space. The original formulation was further improved by Diffusion Transformers (DiTs) (Peebles and Xie, 2023), which demonstrated that standard Transformers provide a highly scalable architecture for the denoiser network. Latent diffusion models can easily incorporate conditioning on additional information like class labels, text prompts, or infilling masks through classifier-based (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2022) as well as finetuning (Zhang et al., 2023; Dai et al., 2023).
However, equivariant networks are computationally expensive and harder to scale than standard Transformers in terms of data and model size. This is especially relevant for diffusion models, where denoisers typically process inputs as fully connected graphs to capture global structure (Joshi, 2020) and are iteratively run hundreds of times during inference. Recent work has challenges the necessity of 3D inductive biases and equivariance for generative structure prediction tasks, showing that standard Transformers can achieve strong performance on biomolecular complexes (Abramson et al., 2024) and small molecule conformations (Wang et al., 2024; O Pinheiro et al., 2023). Non-equivariant models have also shown promising results for protein structure generation (Chu et al., 2024; Martinkus et al., 2024). In the same vein, our work leverages the simplicity and scalability of standard Transformers for generative modelling across both periodic and non-periodic 3D atomic systems, demonstrating that explicit equivariance and molecular inductive biases are not a strict requirement for generating valid and realistic atomic structures at scale.
Appendix B Evaluation Metrics
  • Stability: % of crystals with DFT energy above hull $<$0.0 eV/atom and no. of unique elements >=>=> = 2. (We also report metastability as DFT energy above hull $<$0.1 eV/atom and no. of unique elements >=>=> = 2.)
  • Validity: % of molecules with canonical SMILES string found by RDKit.
  • Uniqueness: % of unique SMILES among valid ones.
Appendix C Additional Results
Extension to MOF generation. Having established that ADiTs benefit from transfer learning to generate high-quality crystals and molecules, we challenged our architecture to generate metal-organic frameworks (MOFs), which represent a more complex class of hybrid materials with metal nodes connected by organic small molecule linkers. We trained ADiTs on an additional 14,000 MOFs of up to 150 atoms from the QMOF database (Rosen et al., 2021) alongside QM9 and MP20, using the same experimental settings and training for 10,000 epochs. We sampled 1,000 MOFs and evaluated their validity using 15 sanity checks from MOFChecker (Jablonka, 2023), including tests for: presence of metal/carbon/hydrogen atoms, atomic overlaps, over/undervalent carbons and nitrogens, missing hydrogens, and excessive partial charges.
Table 4 shows that QMOF-only trained ADiT achieves a 15% overall validity rate for MOF generation, which decreases to 10% with joint training on molecules, crystals, and MOFs simultaneously. However, the jointly trained model takes significant time to train and did not fully converge, suggesting that further improvements in MOF generation may be possible with larger models trained for longer. Comparing validity rates for joint vs. dataset-specific ADiTs shows that the joint model benefits from transfer learning across molecules, crystals, and MOFs, achieving high validity rates earlier in training. Notably, the joint model achieves high validity rates of 91% for crystals and 95% for molecules, matching our best dataset-specific models and being able to additionally generate MOFs. While not directly comparable, models specialized for MOF such as MOFDiff (Fu et al., 2024) are trained on large synthetic dataset of 300,000 MOFs and achieve 30% validity rates based on MOFChecker after DFT relaxation of generated MOFs. Our results are reported without DFT relaxation.